OpenStudy (skullpatrol):

Prove: $$c=\frac{c}{1}$$

OpenStudy (anonymous):

Are you serious?

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

Always True

OpenStudy (anonymous):

c=1/1c

OpenStudy (anonymous):

Not a Quadratic

OpenStudy (isaiah.feynman):

Could use limits.. perhaps.

OpenStudy (isaiah.feynman):

@skullpatrol

OpenStudy (anonymous):

Multiply both sides by \(1\):\[ 1c = c \]Is this sufficient?

OpenStudy (anonymous):

You seem to be assuming the statement is true in order to prove it is true.

OpenStudy (anonymous):

Like this:\[ \lim_{x\to 1}x = 1=-1 \]

OpenStudy (the_fizicx99):

Wouldn't that just be c = c ?

OpenStudy (skullpatrol):

$$\frac{a}{1}=a*\frac{1}{1}$$ because $$\frac{a}{b}=a*\frac{1}{b}$$. $$1*\frac{1}{1}=1$$ because $$a*\frac{1}{a}=1$$. $$1*1=1$$ because $$a*1=a$$. Comparing the two equations above: $$\frac{1}{1}=1$$. $$\therefore \frac{a}{1}=a*\frac{1}{1}=a*1=a$$.

OpenStudy (the_fizicx99):

Exactly

OpenStudy (anonymous):

simple absract algebra ! just define the operation "multiple" and define the identity "1"

OpenStudy (turingtest):

Indeed, multiplicative identity could be made to be zero in the right vector space, for instance.